OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2). A) y = -one thirdx - 6 B) y = -one thirdx - 14 C) y = 3x - 6 D) y = 3x - 14

OpenStudy (godlovesme):

First step can u change 3x-y=-3 into slope-intercept form?

OpenStudy (anonymous):

m =3 @Godlovesme

OpenStudy (anonymous):

@Godlovesme hello? i found M = 3

OpenStudy (godlovesme):

Sorry sorry :/ im bak now

OpenStudy (godlovesme):

3 u mean the slope?

OpenStudy (anonymous):

i guess XD

OpenStudy (godlovesme):

Hmm that wasn't what u meant huh ;P

OpenStudy (godlovesme):

Can u solve for y in the given equation? (:

OpenStudy (anonymous):

does this help XD http://prntscr.com/6x10kc

OpenStudy (anonymous):

that shows the slope :)

OpenStudy (godlovesme):

..online graphing calculator ._. .. unfortunately that doesn't help :P Any hoo lets solve it together.. So first step would be subtracting 3x from both sides

OpenStudy (anonymous):

oh sure I know 3x - y = -3 3x /3 -y = -3 /3 x - y = -1 like that? :)

OpenStudy (godlovesme):

Um. Not quite .. but kinda close (: http://prntscr.com/6x1327 Look how I did it c:

OpenStudy (godlovesme):

Let me know if there is anything confusing (:

OpenStudy (anonymous):

ok :)

OpenStudy (godlovesme):

So do u understand how I solver for y?

OpenStudy (anonymous):

yea makes better sense now XD :D

OpenStudy (godlovesme):

Awesome! :D So if the y-intercept for is Y=mx+b And M stands for slope B stands for y-intercept What is the slope of the above equation?

OpenStudy (godlovesme):

Formula**

OpenStudy (godlovesme):

http://www.algebra-class.com/images/form.gif ^^^ that's what im trying to say :P

OpenStudy (anonymous):

so we have y = 3x + 3 wouldn't I take 3x and 3 subtract 3 from the minus 3x? XD

OpenStudy (godlovesme):

Hmm not really look at the pic and u will understand what I'm asking u for :3

OpenStudy (anonymous):

ok :)

OpenStudy (godlovesme):

@Phoenixbirrd so whats the slope of the equation Y=3x+3

OpenStudy (godlovesme):

@Phoenixbirrd tag me when u get back (:

OpenStudy (anonymous):

i am working it out :D

OpenStudy (godlovesme):

Great but It's obvious hold on one sec

OpenStudy (anonymous):

m=3,b=3 where m is the slope and b is the y-intercept ?? right? :D

OpenStudy (godlovesme):

YESSSSSS!!! :DDDDD AWESOME job! :D

OpenStudy (anonymous):

yay :) so am i done whats the answer now? XD abcd? XD @Godlovesme

OpenStudy (godlovesme):

Yea u got it ABCD ;D Lol not done yet :/ but it gets easier :)

OpenStudy (godlovesme):

So do u know anything about the slope of parallel lines?

OpenStudy (godlovesme):

@Phoenixbirrd tag me when u get bak

OpenStudy (anonymous):

a little XD @Godlovesme

OpenStudy (godlovesme):

What do u know? *hint* it has something to do with their slope :P

OpenStudy (anonymous):

nvm I found it its D thanks bro :) you get a medal :

OpenStudy (godlovesme):

Thanks and good job! :D

OpenStudy (godlovesme):

P.s I'm not bro :P lol